Shift report 20030119 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Jan 19 2003 - 23:22:41 EST

  • Next message: Pawel Staszel: "Re: Shift report 20030119 08:00-16:00"
    ------------------------------------------------------------
    Date: Sun Jan 19 23:22:41 2003, Shift: 16:00-24:00
    Supervisor: JJG., with: 
    Log book#13 pages 96 - 101
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam at start of shift until 18.00. About 80 evts/sek, Inel rate: 4K, H1 :25K, H2: 50K, TOFW 50K, BB: 1K.
    
    2. Many DAQ problems. DAQ kept crashing ev. 15 min. with " spool output stream is missing", "IOServer and RunServer failures". FB 3 and T3 LV crashed repeatedly. 
    Finally "FB4 is not communicating to EVB" and scalers 1,3,5,8,9,11 went to ZERO. I still have not understood why. Is all the vital stuff in FB4? I suspect not and cannot see this in the documentation.
    Reboot FB 4. 
    Beam lost before I could make sure it is all working again.
    Next shift will find out!!
    
    3. The last 48 hrs have been afflicted by many DAQ problems, FB failures etc. Together with Beam instability this does not lead to much opportunity to debug dets. and study timing.
    
    4. At quench all DC cathodes tripped and all TPC drift Voltages. Gas detector anodes are left in 'parked' mode.
    
    5. Decided to move FS to 12  deg to study rates and background there and possible influence on DAQ performance.
    
    
    6. Have accumulated about 850 k trig6 at 4 deg (1/2 B) and about 1 M trig 6 at 4 deg (1/2 A). Probably most of these data have intermittent DC problems along the way (RO failures).
    
    7. Someone should  really fix the missing 'power on' LED's in the FEE hut crates. It is not easy to see if there is a real malfunction when there is a problem with the data.
    
    8. The first 16 channels in H1 are still bad. So it was app. not the cable. IGB does not believe it is software as he rightly points out that the ADC spectra look OK and similar.
    Trigger board? BOB?
    
    9. The deck on the way to the FEE hut is VERY ICY. I almost broke a leg. Should SALT. sooner or later someone will fall -before spring comes.
    
    10. At end of shift still no beam. Due to improbable coinc. of beam and functioning DAQ, I still need the 2 hours to investigate the FS trigger timing and do the relevant supermon plots.
    
    11. Friends, this is my last shift here for a while. 
    I will continue to bug you from NBI (based on first analysis, Dana !).  
    In spite of irritating moments, mostly due to RHIC, I find that large parts of the exp. look in pretty good shape. I am sure that once RHIC settles down - until jan 31 when the water bills can no longer be paid - the final problems can be resolved.
    I wish you all a good and smooth running. 
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Jan 19 2003 - 23:23:14 EST